OpenStudy (anonymous):

Simplify the product using the distributive property. (-4h +2)(3h +7) A. -12h2 - 22h + 14 B. -12h2 - 34h - 14 C. -12h2 + 34h - 14 D. -12h2 + 22h + 14

OpenStudy (anonymous):

@eliassaab

Parth (parthkohli):

\[(-4h + 2)(3h + 7) \implies -4h(3h + 7) + 2(3h + 7) \]

OpenStudy (anonymous):

well i just dont know =(

Parth (parthkohli):

Hmm. Use the distributive property.

OpenStudy (anonymous):

(-4h +2)(3h +7)=-20h

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

F.O.I.L. First. Outer. Inner. Last.

OpenStudy (anonymous):

\[(a+b)(c+d)=ac+ad+bc+bd.\]

OpenStudy (anonymous):

You can immediately eliminate choices B and C because (+2)(+7)=+14.

OpenStudy (anonymous):

so it is A or D

OpenStudy (anonymous):

same thing as foil